The state​ lottery's million-dollar payout provides for ​$1.4 million to be paid in 20 installments of ​$70 comma 000 per payment. The first ​$70, 000 payment is made​ immediately, and the 19 remaining ​$70,000 payments occur at the end of each of the next 19 years. If 9 percent is the discount​ rate, what is the present value of this stream of cash​ flows? If 18 percent is the discount​ rate, what is the present value of the cash​ flows? a. If 9 percent is the discount​ rate, the present value of the annuity due is ​$ nothing. ​(Round to the nearest​ cent.
Answer #1

Present value of annuity due=(1+rate)*Annuity[1-(1+interest rate)^-time period]/rate

1.Present value=1.09*70,000[1-(1.09)^-20]/0.09

=70,000*9.950114779

=$696,508.03(Approx).

2.Present value=1.18*70,000[1-(1.18)^-20]/0.18

=70,000*6.316240867

=$442,136.86(Approx).